Given that 𝐺 = {π‘Ž ∈ ℝ|π‘Ž β‰  βˆ’1} and π‘Ž βˆ— 𝑏 = π‘Ž + 𝑏 + π‘Žπ‘, show that (𝐺, βˆ—) is indeed a group.

1
Expert's answer
2021-11-26T16:28:03-0500

group properties:

associativity:

"(a*b)*c=(\ud835\udc4e + \ud835\udc4f + \ud835\udc4e\ud835\udc4f)*c=\ud835\udc4e + \ud835\udc4f + \ud835\udc4e\ud835\udc4f+c+(\ud835\udc4e + \ud835\udc4f + \ud835\udc4e\ud835\udc4f)c="

"=\ud835\udc4e + \ud835\udc4f + \ud835\udc4e\ud835\udc4f+c+ac+bc+abc"

"a*(b*c)=a*(b+c+bc)=a+b+c+bc+a (b+c+bc)="

"=a+b+c+bc+ab+ac+abc"

"(a*b)*c=a*(b*c)"


unit element e:

"e*a=e+a+ae=a"

"e=0"


inverse element b=a-1:

"a*b=a+b+ab=e=0"

"b=-\frac{a}{a+1}"


so, (𝐺, βˆ—) is a group